# The Impact of Meteorological Factors on Crop Price Volatility in India: Case studies of Soybean and Brinjal

### v1

@@ Abstract
The price volatility of agricultural crops (commodities) is influenced by meteorological variables such as temperature and precipitation (and many other environmental, social and governance factors), which is a critical challenge in sustainable finance, agricultural planning, and policy-making. This paper studies the impact of meteorological variables on price volatility of agricultural crops. As case studies, we choose the two Indian states of Madhya Pradesh (for Soybean) and Odisha (for Brinjal). We employ an Exponential Generalized Autoregressive Conditional Heteroskedasticity (EGARCH) model to estimate the conditional volatility of the log returns of crop prices from 2012 to 2024. This study further explores the cross-correlations between volatility and the meteorological variables. Further, a Granger-causality test is carried out to analyze the causal effect of meteorological variables on the price volatility. Finally, the Seasonal Auto-Regressive Integrated Moving Average with Exogenous Regressors (SARIMAX) and Long Short-Term Memory (LSTM) models are implemented as simple machine learning models of price volatility with meteorological factors as exogenous variables. We believe that this will illustrate the usefulness of simple machine learning models in agricultural finance, and help the farmers to make informed decisions by considering climate patterns and making beneficial decisions with regard to crop rotation or allocations. In general, implementing meteorological factors to assess agricultural performance could help to understand and reduce price volatility and possibly lead to economic stability.
@@ 1 Introduction
A system, natural or man-made, consisting of many components that interact nonlinearly among themselves, is termed a complex system [1, 2]. Complex systems like climate, hydrology, financial markets, etc., exhibit extreme events that have severe impacts on different sectors [3, 4, 5]. Understanding various complex systems is of utmost importance, as these systems and their interactions influence natural, economic, and social environments. Climate is one such complex system that has significant impacts on various fields such as agriculture, human health, and disaster management [6]. Even a slight change in the climatic behavior can have significant impacts; hence, it is essential to analyze the dynamics and influence of climate on various sectors.
Agriculture remains one of the sectors most vulnerable to climate variability, with farmers heavily reliant on meteorological variables like temperature and precipitation for crop production. However, unpredictable fluctuations in these factors pose substantial risks, disrupting planting and harvesting schedules, reducing yields, and leading to unexpected shortages or surpluses in agricultural output [7, 8, 9, 10]. These climate-induced disruptions can have profound consequences for food security and income stability, particularly in regions where agriculture is a primary source of livelihood, such as developing countries [11, 12, 9].
The fluctuations in crop supply directly influence market prices, leading to increased price volatility that can destabilize local economies and threaten the financial well-being of smallholder farmers[13, 9, 10]. Extreme precipitation makes economic inequality worse in nations that rely heavily on agriculture. When precipitation rises by 1.5 standard deviations, low-income groups are disproportionately affected—35 times more so than in nations with less reliance on agriculture. Given this pattern, it is probable that wealth inequality will increase in agricultural areas, especially in Africa, calling for immediate attention to how climate change affects vulnerable groups[14]. Furthermore, the movement of nitrogen from land to aquatic systems is being disrupted by climate change, as rising temperatures in many parts of the United States may offset the benefits of greater precipitation on nitrogen runoff. The intricate relationship between environmental sustainability, economic inequality, and climate change is made clear by these revelations[15]. Long-term climate trends have reduced wheat and barley yields across Europe by 2.5% and 3.8%, with Mediterranean areas facing the most adverse effects. This suggests climate factors explain around 10% of yield stagnation, with policy and agricultural changes likely accounting for the rest[16]. By the end of the century, if global temperature rises above the ideal thresholds for crops (29°C for corn, 30°C for soybeans, and 32°C for cotton), the U.S., which produces 41% of the world’s maize and 38% of its soybeans, may experience production declines of up to 82%[17]. The development of heat-tolerant crops is vital since warming of 2°C and 4°C would likely increase the variability of maize yields, threatening food security and the stability of the grain trade, particularly for the 800 million people living in extreme poverty[18]. Similarly, India’s agriculture sector must adapt its cropping patterns for sustainable growth, particularly after COVID-19 intensified reliance on agriculture for livelihoods. In West Bengal, crop diversification, irrigation, fertilizer, and market access are driving transitions toward non-foodgrains, with climate factors like humidity and temperature playing a role. Long-term sustainability will require enhanced infrastructure, agricultural education, and decentralized support to address local dynamics effectively[19].
Price volatility in agricultural markets has been extensively studied in relation to market forces, but recent studies have increasingly emphasized the role of meteorological variables. Volatility models such as Generalized Autoregressive Conditional Heteroskedasticity (GARCH) and Exponential Generalized Autoregressive Conditional Heteroskedasticity (EGARCH) have been used to analyze time series data, accounting for asymmetry in shocks affecting price returns. Research indicates that weather conditions, especially extreme weather events, significantly impact agricultural productivity and pricing. These conditions lead to fluctuations in supply, thereby affecting prices. For instance, studies utilizing GARCH-MIDAS frameworks have demonstrated that temperature anomalies and precipitation patterns are pivotal in understanding food price volatility. These models reveal that climate change can induce significant volatility in food prices, emphasizing the need for integrating meteorological variables into volatility modeling [13, 9, 10]. Forecasting price volatility through meteorological data offers a strategic approach to managing the risks posed by climate variability. By leveraging predictive models that incorporate meteorological variables, policymakers and financial institutions can develop more effective crop insurance schemes and financial tools to protect farmers from the adverse impacts of climate fluctuations [20, 8, 12]. Empirical results indicate that hybrid ARIMAX-LSTM models outperform traditional methods, achieving lower mean absolute percentage error (MAPE) values, thereby enhancing the accuracy of price forecasts and risk mitigation strategies[13, 10].
This paper focuses on estimating the price volatilities of two critical agricultural commodities in India — Soybean in Madhya Pradesh, and Brinjal in Odisha using the EGARCH model. We selected these crops as they represent two different agriculture markets; Soybean being an export oriented oil seed and Brinjal bring a vegetable which is used for domestic use. Additionally, Soybean is a kharif crop, which is strictly sown during the onset of monsoon (June/July) and harvested during (October) [21, 22]. On the other hand, Brinjal is cultivated all year round without seasonal cycles [23, 24]. Further, we implement SARIMAX and LSTM models to forecast price volatility using the meteorological variables like maximum temperature and precipitation. This approach will help anticipate risk and offer insights into how tailored financial models can mitigate the socio-economic challenges posed by climate-driven agricultural volatility. The findings of the work will provide the following revelation: firstly, the influence of meteorological variables on the commodity prices, and secondly, which model proves a better fit for such analysis.
@@ 2.1 Description
The Government of India’s Directorate of Marketing and Inspection operates the AGMARKNET web portal, which plays a crucial role in disseminating agricultural market information, including arrivals and prices of various agricultural commodities across India. The portal collects data from local agricultural markets through a specially designed application called "Agmark," enabling farmers, traders, and researchers to access real-time market trends and insights. The website for accessing the corresponding data is: AGMARKNET. By providing reliable and transparent information, AGMARKNET aims to support informed decision-making among stakeholders in the agricultural sector, ultimately fostering improved market practices and facilitating effective price discovery. We have taken monthly data from eighteen districts in Madhya Pradesh and six districts in Odisha, covering the period from Jan-2012 to Oct-2024. The district were considered based on the data availability for this period.
@@
The monthly data for precipitation and maximum temperature, measured at 2 meters above ground level from Jan-2012, to Oct-2024, corresponding to the crop price data for Brinjal in Odisha and Soybean in Madhya Pradesh, was sourced from NASA’s POWER Project (Prediction Of Worldwide Energy Resources) and the website for the corresponding the data is NASAPOWER. This data service provides global meteorological and solar information to support applications in renewable energy, sustainable agriculture, and climate-related projects. Powered by NASA’s Earth science satellite observations and reanalysis products, including MERRA-2 (Modern-Era Retrospective analysis for Research and Applications) and CERES (Clouds and the Earth’s Radiant Energy System), the POWER project offers a comprehensive range of data for various sectors. The project primarily focuses on delivering solar irradiance data to optimize solar energy systems, alongside climate variables like temperature, precipitation, humidity, and wind speed, which are critical for crop modeling and weather monitoring in agriculture. It provides global data at high spatial resolution, with options for daily to annual time steps. This data can be accessed via an easy-to-use web interface or programmatically through an API, enabling customized dataset retrieval for specific locations and timeframes.
2.2 Preprocessing, exploratory analyses and visualization
Refer to captionSoybean Price (Rs./Quintal)
Refer to captionBrinjal Price (Rs./Quintal)Year
Figure 1: Monthly Price Analysis. The plots represent the monthly price time series for Soybean and Brinjal (in Rs./Quintal) over the period 2012 - 2024, respectively. The plots show soybean price trends in Madhya Pradesh (MP) and brinjal in Odisha, highlighting seasonal peaks and volatility patterns.
In order to carry out state-level analysis, we have taken the daily average of the commodity prices (soybean in Madhya Pradesh and brinjal in Odisha) from different districts. Further, we have calculated the monthly average for each state which is shown in Fig. 1. The figure represents the monthly price (Rs./Quintal) of soybean in Madhya Pradesh (MP) and brinjal in Odisha from 2012 to 2014. The band (shown by a lighter color) is the one standard deviation price band from the price. Subsequently, we have calculated the log-return of the commodity price which is defined as:
$r_{t}=\ln\Big{(}\frac{P_{t}}{P_{t-1}}\Big{)}=\ln(P_{t})-\ln(P_{t-1}).$
where $r_{t}$ is the log return at time $t$, $P_{t}$ is the price of the commodity at time $t$, and $P_{t-1}$ is the price at time $t-1$.
Figure 2 represents the log return of monthly prices for Soybean and Brinjal, respectively. Log return is used to measure the percentage change in prices, offering a clearer view of relative price movements over time, smoothing out seasonal variations. These plots reflect the price dynamics and volatility, aiding in understanding risk and return in the agricultural markets.
@@ 2.2 Preprocessing, exploratory analyses and visualization
Figure 2: Log Return Analysis. The plots represent the monthly log returns of soybean and brinjal prices respectively. The log returns showcase potential market shocks in the past decade.
Figure 3 displays monthly precipitation and maximum temperature data for Madhya Pradesh and Odisha. All four plots exhibit pronounced annual seasonal patterns. The precipitation plots (in blue) show significant spikes corresponding to India’s monsoon season, when rainfall is highest. The maximum temperature plots (in red) demonstrate clear yearly cycles, with peak temperatures exceeding 40°C.
@@
Figure 3: Meteorological Variables: The plots represent the monthly precipitation and maximum temperature from 2012 to 2024 for Madhya Pradesh (MP) and Odisha.
Table 1: Statistical Summary for Soybean and Brinjal Prices and its Log Returns
Parameter Soybean Prices Brinjal Prices
@@
Mean 3875.51 0.0039 2302.55 0.0089
Standard Deviation 1136.18 0.0737 716.79 0.1992
Minimum 2289.26 -0.3373 1128.12 -0.4284
Maximum 8511.76 0.2273 4872.30 0.5752
Skewness - -0.6102 - 0.2302
Kurtosis - 3.7131 - -0.1736
Table 1 shows the statistical summary for soybean and brinjal monthly prices. The statistics is based on a consistent sample size of 154 observations for each crop. The mean price for soybean is Rs. 3875.51, indicating a significantly higher average compared to brinjal prices, which average is Rs. 2302.55. This disparity highlights a relatively greater market value of soybean. In terms of log returns, soybean exhibit a mean of 0.0039, while for brinjal it is 0.0089, suggesting both crops have shown modest increase in value over time. The standard deviation, which measures price variability, reveals that soybean prices are more volatile, with a standard deviation of 1136.18 compared to 716.79 for brinjal. For log returns, the standard deviations are 0.0737 for soybean and 0.1992 for brinjal, indicating that log return for brinjal exhibits greater variability. The price range for soybean spans from 2289.26 to 8511.76, while brinjal prices range from 1128.12 to 4872.30. This broader price range for soybean suggests a more diverse market response, however the relative change is higher for brinjal. Distribution characteristics further illuminate the behavior of returns. Soybean log returns show a negative skew of -0.6102, indicating a tendency for extreme negative returns, which may signal potential downside risks. In contrast, brinjal log returns exhibit a positive skew of 0.2302, reflecting a slight tendency for extreme positive returns, which could indicate more upside potential. The kurtosis value for soybean log returns is 3.7131, indicating a leptokurtic distribution characterized by a sharper peak and fatter tails, suggesting a higher likelihood of extreme values. Conversely, brinjal log returns have a kurtosis of -0.1736, indicating a flatter distribution with less probability of extreme returns. However, as the value is very close to zero the behavior will be similar to normal distribution.

### v3

@@ Abstract
Climate is an evolving complex system with dynamic interactions and non-linear feedback mechanisms, shaping environmental and socio-economic outcomes. Crop production is highly sensitive to climatic fluctuations (and many other environmental, social and governance factors). This paper studies the price volatility of agricultural crops as influenced by meteorological variables, which is critical for agricultural planning, sustainable finance and policy-making. As case studies, we choose the two Indian states: Madhya Pradesh (for Soybean) and Odisha (for Brinjal/Eggplant). We employ an Exponential Generalized Autoregressive Conditional Heteroskedasticity (EGARCH) model to estimate the conditional volatility of the log returns from 2012 to 2024. We further explore the cross-correlations between price volatility and the meteorological variables followed by a Granger-causal test to analyze the causal effect of meteorological variables on the volatility. The Seasonal Auto-Regressive Integrated Moving Average with Exogenous Regressors (SARIMAX) and Long Short-Term Memory (LSTM) models are implemented as simple machine learning models of price volatility with meteorological factors as exogenous variables. Finally, to capture spatial dependencies in volatility across districts, we extend the analysis using a Conditional Autoregressive (CAR) model to construct monthly volatility surfaces that reflect both local price risk as well as geographic dependence. We believe, this paper will illustrate the usefulness of simple machine learning models in agricultural finance, and help the farmers to make informed decisions by considering climate patterns and making beneficial decisions with regard to crop rotation or allocations. In general, incorporating meteorological factors to assess agricultural performance could help to understand and reduce price volatility and possibly lead to economic stability.
@@ 1 Introduction
Systems–whether natural, like climate and hydrology, or artificial, like financial markets–comprise of numerous independent components whose intricate interactions give rise to emergent behaviours [1, 2, 3]. These behaviours often cannot be predicted by examining individual parts in isolation [4, 5]. Such systems are prone to extreme events with far-reaching consequences across natural, economic, and social environment. Amongst these, climate stands out as an evolving complex system characterized by dynamic interactions and non-linear feedback mechanisms that shape environmental and socio-economic outcomes [6]. Climate change, being one of the most pressing challenges today, exemplifies the multifaceted nature of such systems. India, in particular, ranks sixth among countries most affected by its impact [7]. Given the complexity and feedback-driven interactions characteristic of climate systems, employing advanced analytical approaches is essential for a comprehensive understanding. For instance, dynamical systems theory has been applied to study climate variability in India, revealing significant dynamic shifts during strong El Niño and La Niña events and uncovering larger patterns within climate data[8]. Such methodologies provide deeper insights into the intricate behaviors of climate systems and their broader implications.
Agriculture–especially in developing countries like India–is extremely sensitive to climate fluctuations. Even minor shifts in temperature or precipitation can disrupt planting cycles, reduce yields, and destabilize supply chains, resulting in sharp market fluctuations and increased price volatility [9, 10, 11, 12]. For farmers whose primary source of income is crop yield and sales, agricultural price fluctuations pose severe risks to their livelihood, and threaten food security [13, 14, 11]. In India, where agriculture is both, an important economic sector and a vital source of employment, understanding and predicting the dynamics of crop price volatility is essential. Although traditional research has shown how average weather conditions affect crop yields or overall prices [13], volatility of crop prices due to climate variability has not been explored much.
This study aims to address this gap by applying a complex systems approach to analyze agricultural price volatility. Crop prices are treated as a financial time series to analyze their volatility characteristics, with log returns and squared log returns computed to estimate conditional volatility. Subsequently, the effect of meteorological variables (temperature and precipitation) on crop price volatility is evaluated, using a methodology that allows for the examination of non-linear and asymmetric effects of climate change on agricultural markets. Two distinct crops and regions have been analyzed: soybean in Madhya Pradesh, an export-oriented kharif crop, and brinjal(eggplant) in Odisha, a domestically consumed vegetable grown all year. Soybean in Madhya Pradesh was chosen due to Madhya Pradesh’s dominant role in India’s soybean sector–contributing approximately 42% of national production as of 2023-24[15]. The state, often referred to as the "Soybean State"[15], has the largest area under soybean cultivation, approximately 5.35 million hectares[16]. Brinjal(eggplant) cultivation in Odisha was chosen due to the state’s status as the second-largest producer of the vegetable in India[17]. While West Bengal leads in overall production, its highly diverse climate poses challenges for consistent agricultural studies[18]. In contrast, Odisha offers relatively uniform meteorological conditions across the state, making it a more suitable and controlled environment for focused analysis. A combination of econometric (EGARCH) and machine learning models (SARIMAX and LSTM) is employed to investigate whether and how precipitation and maximum temperature affect crop price volatility, and whether these effects can be accurately forecasted. Further, to enrich the spatial understanding of agricultural price volatility, a Conditional Autoregressive (CAR) model has been applied to construct monthly volatility surfaces across Madhya Pradesh and Odisha. This spatially explicit modelling provides a granular view of volatility distribution, offering a valuable tool for assessing local-level price risk. By integrating high-resolution meteorological data with market prices of the two crops, the study seeks to anticipate risk and offer insights into how tailored financial models can mitigate the socio-economic challenges posed by climate-driven agricultural volatility. The findings of the study highlight the extent to which meteorological variables influence agricultural commodity prices. This analysis aims to elucidate how climate variability translates into price volatility offering insights that are both theoretically and practically relevant for policy-making and risk management.
2 Literature Review
The fluctuations in crop supply directly influence market prices, leading to increased price volatility that can destabilize local economies and threaten the financial well-being of smallholder farmers[19, 11, 12]. Extreme weather events, like precipitation anomalies, have a pronounced impact on economic inequality in nations that rely heavily on agriculture. Research shows that in these nations a 1.5 standard deviation increase in precipitation relative to average levels affects the low-income groups 35 times more in countries where agricultural employment is high (for example, Africa), compared to those with minimal agricultural reliance. This shows that anomalies in rainfall and the level of agricultural dependence are important factors to consider while evaluating the negative impacts of climate change on vulnerable groups[20]. Climate change also disrupts broader environmental systems. For instance, changes in temperature can offset the expected benefits of increased rainfall on nitrogen runoff. Additionally, it also disrupts the movement of nitrogen from land to aquatic systems. The intricate relationship between environmental sustainability, economic inequality, and climate change is made clear by these revelations[21].
Looking ahead, climate projections suggest that crop yields could be reduced to a large extent if global temperatures rise beyond ideal thresholds for crops–29°C for corn, 30°C for soybean, and 32°C for cotton. The U.S., which produces 41% of global maize and 38% of soybean, may experience production declines of up to 82%[22]. The development of heat-tolerant crops is therefore vital since warming of 2°C and 4°C would likely increase the variability of maize yields, threatening food security and the stability of the grain trade, particularly for the 800 million people living in extreme poverty[23]. In West Bengal for instance, transitions toward non-food grains are being influenced by temperature and humidity, pointing to the complex and localized impact of climate change[24]. Thus, India’s agricultural sector must adapt its cropping patterns for sustainable growth.
Although agricultural price volatility has historically been studied through market-centric frameworks, there is a growing recognition of the influence of meteorological drivers. Before exploring advancements in analytic tools and methodologies, it is important to first revisit the classical economic theories that have shaped our understanding of agricultural price volatility. In the early eighties, under the influence of the "Washington Consensus", trade liberalization, reduced state intervention, and reliance on international markets for food security was advocated for[25]. The theoretical foundation of such policies were rooted in welfare economics, which states that free markets guarantee maximum welfare outcomes given economic agents are rational and perfect information exists[26, 27]. Traditional economic models often attribute agricultural price volatility to exogenous shocks affecting supply, such as climate change or pests, coupled with inelasticity of demand. These shocks are typically represented as Gaussian disturbances in applied models[25]. The recommended solution to overcome price instability is trade and private storage. This is because operators are incentivised to buy when prices are low and sell when prices are high, leading to market enlargement and stability through the "law of large numbers"[28, 29]. However, these approaches failed to predict the global food price crisis of 2006-2008 and 2010-2011, raising questions about the understanding of commodity price determinants and the validity of the "rational expectations hypothesis"[25]. Literature demonstrates that markets do not always converge to an equilibrium. This highlights the role of endogenous factors and market imperfections while understanding agricultural price volatility. Newer models such as Generalized Autoregressive Conditional Heteroskedasticity (GARCH)[30] and Exponential Generalized Autoregressive Conditional Heteroskedasticity (EGARCH)[31] have been used to analyze time series data. These models capture asymmetric responses to shocks, making them particularly suitable for agricultural contexts, where weather-induced price shifts can be sharp and skewed. Research employing GARCH-MIDAS models have revealed that temperature anomalies and irregular precipitation are crucial determinants of food price volatility[19, 11, 12]. These insights underscore the need to integrate meteorological variables into volatility modelling. Climate-induced volatility, when appropriately modeled, enables more effective agricultural risk management strategies.
Predictive models that incorporate meteorological variables offer a strategic response to risks posed by climate vulnerability. By embedding temperature and precipitation data, forecasting tools can assist policymakers and financial institutions in designing resilient crop insurance schemes and other mitigation mechanisms[32, 10, 14]. Among various models, hybrid approaches such as ARIMAX-LSTM models outperform traditional methods, achieving lower mean absolute percentage error (MAPE) values thereby enhancing the accuracy of price forecasts and risk mitigation strategies[19, 12]. They capture both linear and non-linear dependencies between climate variables and price volatility. A complex systems approach, as utilised in this study, addresses limitations of traditional linear models by framing agricultural markets as systems with interacting parts–farmers, markets, weather patterns, and policy mechanisms–whose collective behaviour exhibits emergent properties.
This study contributes to the literature by applying these hybrid methodologies to two key agricultural commodities in India: soybean in Madhya Pradesh and brinjal(eggplant) in Odisha. These represent contrasting market contexts– soybean as an export-oriented kharif oilseed, and brinjal(eggplant) as a domestically consumed, non-seasonal vegetable [33, 34, 35, 36]. This allows for a comparative analysis of globally traded versus locally consumed commodities. The EGARCH model is used to estimate volatility, SARIMAX and LSTM models are employed to forecast price dynamics by incorporating meteorological variables like maximum temperature and precipitation. Tersely, this paper contributes to the evolution of agricultural economics by shifting the analytical lens from static, exogenous models, to dynamic climate-sensitive frameworks grounded in complex adaptive systems theory.
@@ 3.1 Description
The Government of India’s Directorate of Marketing and Inspection operates the AGMARKNET web portal, which plays a crucial role in disseminating agricultural market information, including arrivals and prices of various agricultural commodities across India. The portal collects data from local agricultural markets through a specially designed application called "Agmark," enabling farmers, traders, and researchers to access real-time market trends and insights. The website for accessing the corresponding data is: AGMARKNET[37]. By providing reliable and transparent information, AGMARKNET aims to support informed decision-making among stakeholders in the agricultural sector, ultimately fostering improved market practices and facilitating effective price discovery. We have taken monthly data from eighteen districts in Madhya Pradesh and six districts in Odisha, covering the period from Jan-2012 to Oct-2024. The districts were considered based on the data availability for this period.
@@
The monthly data for precipitation and maximum temperature, measured at 2 meters above ground level from Jan-2012, to Oct-2024, corresponding to the crop price data for brinjal(eggplant) in Odisha and soybean in Madhya Pradesh, was sourced from NASA’s POWER Project (Prediction Of Worldwide Energy Resources) and the website for the corresponding the data is NASAPOWER[38]. This data service provides global meteorological and solar information to support applications in renewable energy, sustainable agriculture, and climate-related projects. Powered by NASA’s Earth science satellite observations and reanalysis products, including MERRA-2 (Modern-Era Retrospective analysis for Research and Applications) and CERES (Clouds and the Earth’s Radiant Energy System), the POWER project offers a comprehensive range of data for various sectors. The project primarily focuses on delivering solar irradiance data to optimize solar energy systems, alongside climate variables like temperature, precipitation, humidity, and wind speed, which are critical for crop modelling and weather monitoring in agriculture. It provides global data at high spatial resolution, with options for daily to annual time steps. This data can be accessed via an easy-to-use web interface or programmatically through an API, enabling customized dataset retrieval for specific locations and time frames.
3.2 Preprocessing, Exploratory Analysis and Visualization
To perform a state-level analysis, the monthly closing prices for each district were averaged to derive a single representative price time series for soybean in Madhya Pradesh and brinjal in Odisha. The resulting monthly prices are shown in Fig. 1.
Refer to captionPrice (INR/Quintal)
Refer to captionPrice (INR/Quintal)Year
Figure 1: Monthly Closing Price. The brown and purple lines represent the monthly closing price of soybean and brinjal from 2012 to 2024 respectively while the shaded regions indicate the confidence intervals, calculated as ±$\pm 2$ standard deviation across districts, capturing the variability in prices within each state. (1 Quintal $\equiv$ 100 kg).
The monthly closing prices for soybean show a relatively stable trend from 2012-2019, with prices largely ranging between 3000 and 4000 INR/Quintal. A significant price surge can be observed from 2020 onward peaking sharply in 2021 and early 2022, rising almost above 8000 INR/quintal. This most likely reflects the global market disruptions caused by the COVID-19 pandemic. Post 2022, prices gradually decline and show a stable pattern around 4000 INR/Quintal. Brinjal prices on the other hand, display a more volatile and erratic pattern. Prices generally range between 1000-3000 INR/Quintal from 2012-2019. Post this, an upward shift can be seen, where prices range between 4000-5000 INR/Quintal. The plot reflects the sensitivity of brinjal to localized factors like perishability, weather events, or limited post-harvest infrastructure, making its prices regionally dependent.
Following this, the log returns of the crop prices is computed as:
$r_{t}=\ln\Bigg{(}\frac{P_{t}}{P_{t-1}}\Bigg{)}=\ln(P_{t})-\ln(P_{t-1}),$
where $r_{t}$ denotes the log returns at time $t$, while $P_{t}$ and $P_{t-1}$ represent the agricultural commodity prices at times $t$ and $t-1$, respectively. Figure 2 represents the log returns of monthly prices for soybean and brinjal respectively, measuring relative price changes over time while normalising variance. It is particularly useful for estimating volatility and measuring risk in commodity markets.
@@ 3.2 Preprocessing, Exploratory Analysis and Visualization
Figure 2: Monthly Log Returns. The blue and purple lines represent the monthly log returns of soybean and brinjal prices respectively.
Log returns highlight potential market shocks over time. In the above figure, log returns for soybean show relatively low and stable price fluctuations with a few spikes–around 2012-2013 and early 2022–indicating brief but sharp changes in prices likely driven by external shocks. In contrast, log returns of brinjal prices show frequent and larger fluctuations, reflecting higher price volatility. These patterns suggest that brinjal prices are more unstable and reactive to short-term disturbances compared to soybean, underlying the greater market risk associated with perishable commodities.
Figure 3 represents the squared log returns of the monthly prices for soybean and brinjal respectively. The squared log returns exaggerate large spikes in price changes, making periods of high volatility more visually prominent. The plots show that the price volatility of brinjal is larger than soybean due to idiosyncratic factors such as localized supply-demand fluctuations and market structure.
Refer to captionSquared Log Returns
Refer to captionSquared Log ReturnsYear
Figure 3: Monthly Squared Log Returns. The brown and purple lines represent the monthly squared log returns of soybean and brinjal prices respectively.
Figure 4 displays monthly precipitation and maximum temperature for Madhya Pradesh and Odisha. All four plots exhibit pronounced annual seasonal patterns. The precipitation chart (in blue) show significant spikes corresponding to India’s monsoon season, when rainfall is highest. The maximum temperature plots (in red) demonstrate clear yearly cycles, with peak temperatures exceeding 40°C.
@@
Figure 4: Meteorological Variables. The blue and red lines represent the monthly precipitation and maximum temperature from 2012 to 2024 respectively for Madhya Pradesh (left column) and Odisha (right column).
Table 1 shows the statistical summary for soybean and brinjal monthly prices. Interestingly, soybean log returns show a negative skew of -0.9839, indicating a tendency for extreme negative returns, which may signal potential downside risks. In contrast, brinjal log returns exhibit a positive skew of 0.3889, reflecting a slight tendency for extreme positive returns, which could indicate more upside potential. The kurtosis value for soybean log returns is 8.0946, indicating a leptokurtic distribution characterized by a sharper peak and fatter tails, suggesting a higher likelihood of extreme values. Conversely, brinjal log returns have a kurtosis of 0.2779, indicating a flatter distribution with less probability of extreme returns.
Table 1: Statistical Summary for Soybean and Brinjal Prices and their Log Returns
Parameter Soybean Brinjal
@@
Mean 3867.74 0.0040 2285.16 0.0077
Standard Deviation 1150.03 0.0905 744.22 0.2736
Minimum 2269.71 -0.5164 1045.83 -0.7154
Maximum 8511.76 0.3115 4250.00 0.7862
Skewness - -0.9839 - 0.3889
Kurtosis - 8.0946 - 0.2779